OpenStudy (anonymous):

Wanda opened a savings account 20 years ago with a deposit of $3,150.37. The account has an interest rate of 4.1% compounded daily. How much interest has Wanda earned?

OpenStudy (anonymous):

Use the interest formula: I=PRT

OpenStudy (anonymous):

No, compound interest formula!

OpenStudy (anonymous):

Okay, M = P( 1 + i )n M is the final amount including the principal. P is the principal amount. i is the rate of interest per year. n is the number of years invested.

OpenStudy (anonymous):

Pay attention at the key word "Compounded daily"

OpenStudy (anonymous):

Now plug, it in, @Sebivid , and tell us.

OpenStudy (anonymous):

thats the wrong formula

OpenStudy (anonymous):

I know i have to use the Compound interest formula but i dont know how to slove for I

OpenStudy (anonymous):

Put the compound interest formula here, and we can see.

OpenStudy (anonymous):

Compounded daily: i = r/ 365 = .041/365 = ... t = 20 yrs * 365 days = ...

OpenStudy (anonymous):

\[A=P(1+r/n) ^{nt}\]

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

$7,152.59 Just got it!! thanks for the help

OpenStudy (anonymous):

Not yet, in order to find interest: I = A - P

OpenStudy (anonymous):

Ohh right

OpenStudy (anonymous):

4002.22 Thank You again!!

OpenStudy (anonymous):

:)